# Investigating symptom duration using current status data: a case study of post-acute COVID-19 syndrome

### v2

@@ Acknowledgements
We refer to $\tau_{0}$ as the scale factor. Construction of confidence intervals requires estimation of $\tau_{0}$. We note that both $\kappa_{0}$ and $f_{0}$ can be estimated by simply plugging in available estimates of $\mu_{0}$ and $g_{0}$, with the empirical distribution of $W_{1},W_{2},\ldots,W_{n}$ used to estimate $Q_{0}$. To estimate $\theta_{0}^{\prime}(y_{0})$, the derivative of $y\mapsto\theta_{0}(y)$ evaluated at $y_{0}$, there are a number of available options. In our implementation, we construct a monotone Hermite spline approximation of $\theta_{n}$ and perform numerical differentiation.
S2.   Regression analysis
@@
Many popular regression models used in survival analysis have been adapted for use with current status data, including the Cox proportional hazards model [44], the proportional odds model [45], and the accelerated failure time model [46]. These methods are generally valid under the same assumptions on the survey response process as the adaptation of the CIR method described above. They do, however, make additional assumptions on the relationship between $T$ and $W$.
@@
Fitting regression models using current status data is generally more complicated than with right censored data. For the Cox model, unlike in the right-censored setting, estimation of the regression coefficients cannot be completely separated from estimation of the baseline hazard function through the partial likelihood [44]. Nonetheless, estimation and valid inference is still attainable for the regression coefficients, with the main difficulties being computational in nature. Recent work has focused on developing more efficient algorithms for computing estimates of the regression coefficients via maximum likelihood [36]. Calculation of the observed Fisher information, which is used to compute analytic standard error estimates, is generally impractical. The bootstrap has been proposed as an alternative approach [47], although as of yet there is no theoretical proof of its validity in this context. In Section S3, we present results of a simulation study investigating the performance of the nonparametric bootstrap for construction of confidence intervals using current status data for inference on regression coefficients of a Cox model. Wald-type bootstrap confidence intervals based on a normal approximation demonstrate good performance, while percentile bootstrap confidence intervals tend to be slightly anticonservative. Based on these findings, we make use of the Wald-type interval in the regression analyses we report.
@@
As with the Cox model, the other regression models described above do not need to be altered to account for survey nonresponse, under the assumption that the regression model is correctly specified and includes all covariates that inform the response mechanism [32].
S3.   Simulation studies
S3.1   Causal isotonic regression simulations
@@
In this section, we describe a simulation study conducted to evaluate the performance of the proposed extended CIR accounting for survey nonresponse. While we expect the complete case CIR procedure to perform well when follow-up is unlimited, we expect decreased performance when termination of follow-up precludes survey response for some study participants. In contrast, we expect the extended CIR procedure to perform well in both of these scenarios.
@@
In this experiment, the data were generated according to a proportional hazards model. We simulated a covariate vector $W=(W_{1},W_{2},W_{3})$ of three independent components, each distributed according to a Unif({-1,1}) distribution. Given covariate vector $W=w$, we then simulated the survey response time $Y_{*}$ and event time $T$ from independent Weibull distributions with shape parameter 0.75 and scale parameter $\exp(\tfrac{2}{5}w_{1}-\tfrac{1}{5}w_{2})$. For computational purposes, $Y_{*}$ was set to the closest point on a grid of 50 values evenly spaced on the quantile scale. In other words, the values of $Y_{*}$ were coarsened to 50 unique values.
@@
For each simulation replicate, the aim was to estimate the distribution function of $T$ over the interval $[t_{0},t_{1}]$ with $t_{0}=0.02$ (corresponding to the 5th percentile of observed response times) and $t_{1}=1.5$. The maximum follow-up time was set to $c_{0}\in\{2.1,1.8,1.65\}$. We refer to the settings corresponding to these values of $c_{0}$ as Scenarios 1, 2 and 3, respectively. Under these three choices of $c_{0}$, $t_{1}$ corresponded to the 90th, 95th, or 97.5th quantile of observed survey response times, respectively. This allowed us to explore whether performance of the complete case CIR or extended CIR procedures would degrade if $c_{0}$ was chosen such that nonresponse was more prevalent and $t_{1}$ was nearer the edge of the support of observed response times. For each value of $c_{0}$, the observed data consisted of $n$ independent replicates of $(W,Y,\Delta)$ with $Y:=\min(Y_{*},c_{0})$ and $\Delta:=\mathbbm{1}(Y_{*}<c_{0})\mathbbm{1}(T\leq Y)$. We generated datasets of size $n=\{500,1000,1500,2000\}$.
@@
The survival function was estimated either using only complete cases (i.e., survey respondents, or those with $Y_{*}<c_{0}$) with the existing CIR procedure described in [30] or using all participants with the extended CIR procedure proposed here. Nominal 95% confidence intervals were constructed by estimating the scale factor and using the quantiles of the Chernoff distribution. The nuisance function $\mu$ was estimated using Super Learner [33] with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines [48], generalized additive models [49], and random forests [50]. The conditional density $\pi$ was estimated using haldensify [35] with 10-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. For each simulation replicate, the survival function of $T$ was estimated at 100 time-points, evenly spaced on the quantile scale, between $t_{0}=0.02$ and $t_{1}=1.5$. For each setting and method, we conducted 1000 simulation replicates. Pointwise and integrated (over all 100 time-points) performance was quantified using the empirical bias and empirical confidence interval coverage.
@@
From Figures S1 and S2, we see that the extended CIR generally has smaller bias compared to the complete case CIR, although both methods have generally inflated bias near the estimation boundaries. From Figures S3 and S4, we observe that the extended CIR demonstrates strong confidence interval coverage across nonresponse scenarios, while the complete case CIR tends to be somewhat anticonservative on average.
@@
Figure S1: Pointwise bias of the CIR procedure for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from smallest to largest degree of nonresponse. The top and bottom rows denote the standard CIR method using only complete cases and the extended CIR method accounting for nonresponse. The $x$-axis displays the true value of the distribution function. The black line denotes zero bias.
@@
Figure S2: Integrated absolute bias of the CIR procedure for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from lowest to highest nonresponse level. The top and bottom rows denote the standard CIR method using only complete cases and the extended CIR method accounting for nonresponse. The $x$-axis corresponds to sample size. The black line denotes zero bias.
@@
Figure S3: Confidence interval coverage using the CIR procedure for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from smallest to largest degree of nonresponse. The top and bottom rows denote the standard CIR method using only complete cases and the extended CIR method accounting for nonresponse. The $x$-axis displays the true value of the distribution function. The black line denotes the nominal 95% coverage level.

### v3

@@ Acknowledgements
Construction of confidence intervals requires estimation of the scale factor $\tau_{0}(y_{0})$. We note that $\kappa_{0}(y_{0})$ can be estimated by simply plugging in available estimates of $\mu_{0}$ and $g_{0}$, with the empirical distribution of $W_{1},W_{2},\ldots,W_{n}$ used to estimate $Q_{0}$. To estimate $f_{0}$, we note that $f_{0}(y)=\int\pi_{0}(y\,|\,w)Q_{0}(dw)$, which be estimated by $\frac{1}{n}\sum_{i=1}^{n}\pi_{n}(y\,|\,W_{i})$ for a conditional density estimator $\pi_{n}$. To estimate $\theta_{0}^{\prime}(y_{0})$, the derivative of $y\mapsto\theta_{0}(y)$ evaluated at $y_{0}$, there are a number of available options. In our implementation, we construct a monotone Hermite spline approximation of $\theta_{n}$ and perform numerical differentiation.
S2.   Sensitivity analysis
In practice, there may be concern that conditional independence of $T$ and $Y_{*}$ given $W$ does not hold, potentially rendering the proposed estimation procedure systematically biased. In this section, we describe an approach to sensitivity analysis with respect to this key assumption. This sensitivity analysis approach employs copulas, which are commonly used in multivariate analysis to model the dependence between two or more variables.
A copula is a multivariate cumulative distribution function, indexed by an association parameter $\alpha$, that describes the joint distribution of two or more random variables, each with a uniform marginal distribution. Sklar’s theorem (Sklar,, 1959) states that any multivariate distribution can be expressed in terms of its marginal distributions and a copula. Throughout this section, we use $F$ and $f$ to denote generic distribution and density functions, respectively, with a subscript indicating the distribution and random variable corresponding to the given distribution/density function, e.g., $F_{P;T\,|\,W}(t\,|\,w):=P(T\leq t\,|\,W=w)$.
To begin, suppose that for some copula $C_{\alpha}$ the joint distribution function of $T$ and $Y$ given $W=w$ can be expressed as $F_{P;T,Y\,|\,W}(t,y\,|\,w)=C_{\alpha}\{F_{P;T\,|\,W}(t\,|\,w),F_{P;Y\,|\,W}(y\,|\,w)\}$ for all $t$, $y$ and $w$. It is not difficult to show then that $F_{P;T\,|\,Y,W}(t\,|\,y,w)=h_{\alpha}\{F_{P;T\,|\,W}(t\,|\,w),F_{P;Y\,|\,W}(y\,|\,w)\}$, where $h_{\alpha}$ is defined as the mapping $(u,v)\mapsto\left.\frac{\partial}{\partial\tilde{v}}C_{\alpha}(u,\tilde{v})\right|_{\tilde{v}=v}$. Denoting by $m_{\alpha}(u,v)$ the value $m$ such that $h_{\alpha}(m,v)=u$, this yields that $F_{P;T\,|\,W}(t\,|\,w)=m_{\alpha}\{F_{P;T\,|\,Y,W}(t\,|\,y,w),F_{P;Y\,|\,W}(y\,|\,w)\}$, which implies, in particular, that
$F_{P;T\,|\,W}(y\,|\,w)$ $=\ m_{\alpha}\{F_{P;T\,|\,Y,W}(y\,|\,y,w),F_{P;Y\,|\,W}(y\,|\,w)\}$
$=\ m_{\alpha}\{\mu_{P}(y,w),F_{P;Y\,|\,W}(y\,|\,w)\}\ .$
Defining the mapping $\theta_{P,\alpha}:y_{0}\mapsto E_{P}[m_{\alpha}\{\mu_{P}(y_{0},W),F_{P;Y\,|\,W}(y_{0}\,|\,W)\}]$, we find thus that $\theta_{P_{0},\alpha}(y_{0})$ identifies the target parameter $\psi_{0}(y_{0})$ under the dependence copula model considered. Applying the domain transformation $F_{P;Y}$, we obtain the identified primitive function given, for each $\alpha$, by
$\Gamma_{P,\alpha}:y_{0}\mapsto\iint\mathbbm{1}(u\leq y_{0})m_{\alpha}\{\mu_{P}(u,w),F_{P;Y\,|\,W}(u\,|\,w)\}F_{P;Y}(du)F_{P;W}(dw)\ .$
When $\alpha=0$, this yields exactly the identification $\Gamma_{P}$ obtained under the assumption that $Y$ and $T$ are conditionally independent given $W$. Our proposed sensitivity analysis approach entails deriving the efficient influence function of this primitive, constructing a debiased one-step estimator, and using this estimator in the CIR procedure outlined in Section S1.2. In this sense, this sensitivity analysis consists of a strict generalization of the extended CIR procedure developed under conditional independence.
[…]
When $C_{\alpha}$ is taken to be the Frank copula, it is possible to show that
$m_{\alpha}(u,v)$ $=-\frac{1}{\alpha}\log\left\{1-\frac{u(1-e^{-\alpha})}{e^{-\alpha v}+u(1-e^{-\alpha v})}\right\}.$
Here and after, for compactness we revert to the notation used in Section S1, with $\mu_{P}(y,w)=E_{P}(\Delta\,|\,W=w,Y=y)$, $g_{P}(y,w)=\pi_{P}(y\,|\,w)/f_{P}(y)$, $F_{P}$ the marginal distribution function of $Y$, and $Q_{P}$ the marginal distribution function of $W$. Furthermore, we use $H_{P}$ to denote the conditional cumulative distribution function of $Y$ given $W$, that is, $H_{P}(y,w):=P(Y\leq y\,|\,W=w)$. We define $\lambda_{P;\alpha}:y\mapsto\int m_{\alpha}\{\mu_{P}(y,w),F_{P;Y\,|\,W}(y\,|\,w)\}F_{P;W}(dw)$. Then, using the delta method for influence functions and defining
$r_{\alpha}(u,v):=\left\{e^{-\alpha v}+u\left(e^{-\alpha}-e^{-\alpha v}\right)\right\}\left\{e^{-\alpha v}+u\left(1-e^{-\alpha v}\right)\right\},$
we find that the nonparametric efficient influence function $D^{*}_{0,\alpha,y_{0}}$ of $P\mapsto\Gamma_{P,\alpha}(y_{0})$ evaluated at $P_{0}$ is given pointwise by
$(\delta,y,w)\mapsto$ $\frac{\left(1-e^{-\alpha}\right)\mathbbm{1}(y\leq y_{0})\left\{\delta-\mu_{0}(y,w)\right\}e^{-\alpha H_{0}(y,w)}}{\alpha g_{0}(y,w)r_{\alpha}\{\mu_{0}(y,w),H_{0}(y,w)\}}-2\int_{u\leq y_{0}}\lambda_{0;\alpha}(u)F_{0}(du)$
$+\mathbbm{1}(y\leq y_{0})\lambda_{0;\alpha}(y)+\int_{u\leq y_{0}}m_{\alpha}\{\mu_{0}(u,w),H_{0}(u,w)\}F_{0}(du)$
$+\left(1-e^{-\alpha}\right)\int_{u\leq y_{0}}\frac{\left\{\mathbbm{1}(y\leq u)-H_{0}(u,w)\right\}\mu_{0}(u,w)e^{-\alpha H_{0}(u,w)}\left\{1-\mu_{0}(u,w)\right\}}{r_{\alpha}\{\mu_{0}(u,w),H_{0}(u,w)\}}F_{0}(du)\ .$
We note that $\lim_{\alpha\to 0}m_{\alpha}(u,v)=u$, which can be used to verify that the above expression simplifies to the efficient influence function for $P\mapsto\Gamma_{P}(y_{0})$ evaluated at $P_{0}$ as $\alpha\to 0$. Using the empirical distributions $F_{n}$ and $Q_{n}$ to estimate $F_{0}$ and $Q_{0}$, respectively, and with generic estimators $\mu_{n}$, $g_{n}$ and $H_{n}$ of $\mu_{0}$, $g_{0}$ and $H_{0}$, respectively, the one-step estimator takes the form
$\frac{1}{n}\sum_{i=1}^{n}\frac{\mathbbm{1}(Y_{i}\leq y_{0})\left\{\Delta_{i}-\mu_{n}(Y_{i},W_{i})\right\}e^{-\alpha H_{n}(Y_{i},W_{i})}(1-e^{-\alpha})}{\alpha g_{n}(Y_{i},W_{i})r_{\alpha}\{\mu_{n}(Y_{i},W_{i}),H_{n}(Y_{i},W_{i})\}}+\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(Y_{i}\leq y_{0})\lambda_{n,\alpha}(Y_{i})$
$+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\mathbbm{1}(Y_{j}\leq y_{0})\left\{\mathbbm{1}(Y_{i}\leq Y_{j})-H_{n}(Y_{j},W_{i})\right\}\mu_{n}(Y_{j},W_{i})e^{-\alpha H_{0}(Y_{j},W_{i})}\left\{1-\mu_{n}(Y_{j},W_{i})\right\}\left(1-e^{-\alpha}\right)}{r_{\alpha}\{\mu_{n}(Y_{j},W_{i}),H_{n}(Y_{j},W_{i})\}}\ .$
To perform the sensitivity analysis, we use the above one-step estimator in the extended CIR procedure over a range of values of the association parameter $\alpha$.
As noted in the main text, the association parameter $\alpha$ is not identified by the observed data and therefore cannot be estimated. To show this, we construct two distributions $P_{1}$ and $P_{2}$ for the ideal data unit $(T,Y,W)$, with respective association parameters $\alpha_{1}$ and $\alpha_{2}$ with $\alpha_{1}\neq\alpha_{2}$, that imply the same distribution for the observed data unit $(\Delta,Y,W)$. Given a marginal distribution function $F_{W}$ of $W$, we set $F_{P_{1};W}:=F_{W}$ and $F_{P_{2};W}:=F_{W}$. Given a conditional distribution function $F_{Y\,|\,W}$ of $Y$ given $W$, we set $F_{P_{1};Y\,|\,W}:=F_{Y\,|\,W}$ and $F_{P_{2};Y\,|\,W}:=F_{Y\,|\,W}$. Given a conditional distribution function $F_{T\,|\,W}$ of $T$ given $W$, we set $F_{P_{1};T\,|\,W}=F_{T\,|\,W}$ and define $F_{P_{2};T\,|\,W}$ pointwise as
$F_{P_{2};T\,|\,W}(t\,|\,w)=m_{\alpha_{2}}\left[h_{\alpha_{1}}\left\{F_{T\,|\,W}(t\,|\,w),F_{Y\,|\,W}(t\,|\,w)\right\},F_{Y\,|\,W}(t\,|\,w)\right]\ .$
We note then that $P_{1}$ and $P_{2}$ imply, on one hand, that
$P_{\text{obs},1}(\Delta=1\,|\,Y=y,W=w)\ =\ P_{1}(T\leq y\,|\,Y=y,W=w)$
$\hskip 56.9055pt=\ h_{\alpha_{1}}\left\{F_{P_{1};T\,|\,W}(y\,|\,w),F_{P_{1};Y\,|\,W}(y\,|\,w)\right\}$
$\hskip 56.9055pt=\ h_{\alpha_{1}}\left\{F_{T\,|\,W}(y\,|\,w),F_{Y\,|\,W}(y\,|\,w)\right\}\ ,$
and, on the other hand, that
$P_{\text{obs},2}(\Delta=1\,|\,Y=y,W=w)\ =\ P_{2}(T\leq y\,|\,Y=y,W=w)$
$\hskip 56.9055pt=\ h_{\alpha_{2}}\left\{F_{P_{2};T\,|\,W}(y\,|\,w),F_{P_{2};Y\,|\,W}(y\,|\,w)\right\}$
$\hskip 56.9055pt=\ h_{\alpha_{2}}\left(m_{\alpha_{2}}\left[h_{\alpha_{1}}\left\{F_{T\,|\,W}(y\,|\,w),F_{Y\,|\,W}(y\,|\,w)\right\},F_{Y\,|\,W}(y\,|\,w)\right],F_{Y\,|\,W}(y\,|\,w)\right)$
$\hskip 56.9055pt=\ h_{\alpha_{1}}\left\{F_{T\,|\,W}(y\,|\,w),F_{Y\,|\,W}(y\,|\,w)\right\}\ ,$
implying that $P_{\text{obs},1}(\Delta=1\,|\,Y=y,W=w)=P_{\text{obs},2}(\Delta=1\,|\,Y=y,W=w)$. It can similarly be shown that $P_{\text{obs},1}(\Delta=0\,|\,Y=y,W=w)=P_{\text{obs},2}(\Delta=0\,|\,Y=y,W=w)$. Thus, the distribution of $\Delta$ given $(Y,W)$ is the same under $P_{\text{obs},1}$ and $P_{\text{obs},2}$. Of course, by definition, it is also the case that the distribution of $(Y,W)$ is the same under $P_{\text{obs},1}$ and $P_{\text{obs},2}$. Hence, we have exhibited two distributions $P_{1}$ and $P_{2}$ for the full data unit that follow the copula model with distinct indices $\alpha_{1}\neq\alpha_{2}$ but also imply the same distribution for the observed data unit. Thus, the copula parameter is not identifiable from the observed data without any additional assumption.
S3.   Regression analysis
@@
Many popular regression models used in survival analysis have been adapted for use with current status data, including the Cox proportional hazards model (Huang,, 1996), the proportional odds model (Rossini and Tsiatis,, 1996), and the accelerated failure time model (Rabinowitz et al.,, 1995). These methods are generally valid under the same assumptions on the survey response process as the adaptation of the CIR method described above. They do, however, make additional assumptions on the relationship between $T$ and $W$.
@@
Fitting regression models using current status data is generally more complicated than with right censored data. For the Cox model, unlike in the right-censored setting, estimation of the regression coefficients cannot be completely separated from estimation of the baseline hazard function through the partial likelihood (Huang,, 1996). Nonetheless, estimation and valid inference is still attainable for the regression coefficients, with the main difficulties being computational in nature. Recent work has focused on developing more efficient algorithms for computing estimates of the regression coefficients via maximum likelihood (Anderson-Bergman,, 2017). Calculation of the observed Fisher information, which is used to compute analytic standard error estimates, is generally impractical. The bootstrap has been proposed as an alternative approach (Pan,, 1999), although as of yet there is no theoretical proof of its validity in this context. In Section S4, we present results of a simulation study investigating the performance of the nonparametric bootstrap for construction of confidence intervals using current status data for inference on regression coefficients of a Cox model. Wald-type bootstrap confidence intervals based on a normal approximation demonstrate good performance, while percentile bootstrap confidence intervals tend to be slightly anticonservative. Based on these findings, we make use of the Wald-type interval in the regression analyses we report.
@@
As with the Cox model, the other regression models described above do not need to be altered to account for survey nonresponse, under the assumption that the regression model is correctly specified and includes all covariates that inform the response mechanism (Little and Rubin,, 2019).
S4.   Simulation studies
S4.1   Simulations comparing extended CIR to alternative methods
@@
In this section, we describe a simulation study conducted to evaluate the performance of the proposed extended CIR accounting for survey nonresponse. While we expect the complete case CIR procedure to perform well when follow-up is unlimited, we expect decreased performance when termination of follow-up precludes survey response for some study participants. In contrast, we expect the extended CIR procedure to perform well in both of these scenarios. We also evaluated the nonparametric maximum likelihood estimator (NPMLE), which does not account for possible dependence between response times and symptom resolution, computed using only survey respondents.
@@
In this experiment, the data were generated according to a proportional hazards model. We simulated a covariate vector $W=(W_{1},W_{2},W_{3})$ of three independent components, each distributed according to a Unif({-1,1}) distribution. Given covariate vector $W=(w_{1},w_{2},w_{3})$, we then simulated the survey response time $Y_{*}$ and event time $T$ from independent Weibull distributions with shape parameter 0.75 and scale parameters respectively set to
$\exp(\tfrac{2}{5}w_{1}-\tfrac{1}{5}w_{2}+\tfrac{1}{10}w_{3}+\tfrac{1}{10}w_{1}w_{2}+\tfrac{1}{10}w_{1}w_{3}-\tfrac{1}{10}w_{2}w_{3})\ ,$
$\exp(\tfrac{2}{5}w_{1}-\tfrac{1}{5}w_{2}+\tfrac{1}{10}w_{3}+\tfrac{2}{5}w_{1}w_{2}+\tfrac{2}{5}w_{1}w_{3}-\tfrac{2}{5}w_{2}w_{3})\ .$
For computational purposes, $Y_{*}$ was set to the closest point on a grid of 50 values evenly spaced on the quantile scale. In other words, the values of $Y_{*}$ were coarsened to 50 unique values.
@@
For each simulation replicate, the aim was to estimate the distribution function of $T$ over the interval $[t_{0},t_{1}]$ with $t_{0}=0.02$ (corresponding to the 5th percentile of observed response times) and $t_{1}=1.5$. The maximum follow-up time was set to $c_{0}\in\{2.1,1.8,1.65\}$. We refer to the settings corresponding to these values of $c_{0}$ as Scenarios 1, 2 and 3, respectively. Under these three choices of $c_{0}$, $t_{1}$ corresponded to the 90th, 95th, or 97.5th quantile of observed survey response times, respectively. This allowed us to explore whether performance of the complete case CIR or extended CIR procedures would degrade if $c_{0}$ was chosen such that nonresponse was more prevalent and $t_{1}$ was nearer the edge of the support of observed response times. For each value of $c_{0}$, the observed data consisted of $n$ independent replicates of $(W,Y,\Delta)$ with $Y:=\min(Y_{*},c_{0})$ and $\Delta:=\mathbbm{1}(Y_{*}<c_{0})\mathbbm{1}(T\leq Y)$. We generated datasets of size $n\in\{500,1000,1500,2000\}$. For each simulation setting, we performed 500 simulation replicates.
@@
The survival function was estimated (1) using only complete cases (i.e., survey respondents, or those with $Y_{*}<c_{0}$) with the existing CIR procedure described in (Westling et al.,, 2020), (2) using all participants with the extended CIR procedure proposed here, and (3) using the NPMLE, as implemented in the survival software package (Therneau,, 2022), computed only using complete cases. For the two CIR procedures, nominal 95% confidence intervals were constructed by estimating the scale factor and using the quantiles of the Chernoff distribution. The nuisance function $\mu$ was estimated using Super Learner (van der Laan et al.,, 2007) with a library consisting of the marginal mean, logistic regression, multivariate adaptive regression splines (Friedman,, 1991), generalized additive models (GAMs) (Hastie,, 2017), and random forests (Wright and Ziegler,, 2015). The conditional density $\pi$ was estimated using haldensify (Hejazi et al.,, 2022) with 10-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. For each simulation replicate, the survival function of $T$ was estimated at 100 time-points, evenly spaced on the quantile scale, between $t_{0}=0.02$ and $t_{1}=1.5$. For each setting and method, we conducted 1000 simulation replicates. Pointwise and integrated (over all 100 time-points) performance was quantified using the empirical bias and empirical confidence interval coverage.
@@
From Figures S1 and S2, we see that the extended CIR generally has smaller bias compared to the complete case CIR and to the NPMLE, although all methods have generally inflated bias near the estimation boundaries. From Figures S3 and S4, we observe that the extended CIR demonstrates strong confidence interval coverage across nonresponse scenarios, while the complete case CIR tends to be somewhat anticonservative on average, and the NPMLE is highly conservative on average.
@@
Figure S1: Pointwise bias of procedures for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from smallest to largest degree of nonresponse. From top to bottom, the rows denote the standard CIR method using only complete cases, the extended CIR method accounting for nonresponse, and the NPMLE. The $x$-axis displays the true value of the distribution function. The black line denotes zero bias.
@@
Figure S2: Integrated absolute bias of procedures for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from lowest to highest nonresponse level. The three lines in color denote the standard CIR method using only complete cases, the extended CIR method accounting for nonresponse, and the NPMLE. The $x$-axis corresponds to sample size. The black line denotes zero bias.
@@
Figure S3: Confidence interval coverage using the CIR procedure for estimating the survival function under current status sampling with nonresponse. From left to right, the columns represent nonresponse Scenarios 1, 2 and 3, ordered from smallest to largest degree of nonresponse. From top to bottom, the rows denote the standard CIR method using only complete cases, the extended CIR method accounting for nonresponse, and the NPMLE. The $x$-axis displays the true value of the distribution function. The black line denotes the nominal 95% coverage level.